Coca-Cola is carbonated by injection with carbon dioxide gas. Under what conditions is carbon dioxide gas most soluble?
Answer #1

At room temperature the solubility of carbon dioxide is about 90 cm3 of CO2 per 100 mL of water. In aqueous solution, carbon dioxide exists in many forms. First it simply dissolves as

           CO2 (g) = CO2 (aq)

Reaction of CO2 with water produces carbonic acid

               CO2 (aq) + H2O (l) = H2CO3 (aq)

Then it dissociates according to the following equation

             H2CO3 (aq) = H+ (aq) + HCO3- (aq)

             HCO3- (aq) = H+ (aq) + CO32- (aq)

At high pressure the solubility of Carbon Dioxide in water increases many fold.